OpenStudy (anonymous):

Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $21.95 per month plus an additional $0.07 per minute. Cell Best has a plan that costs $25.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month? 0.12m = 4.00 21.95 + 0.07m = 26.00 21.95 + 0.07m = 25.95 + 0.05m 25.95 + 0.05m = 22.02

OpenStudy (anonymous):

21.95+.07m = 25.95 +.05m (subtract .05m from both sides) 21.95 + .02m = 25.95 (subtract 21.95 from both sides) .02m = 4.00 (divide both sides by .02) m = 200 Check both prices with 200 minutes: 21.95 + 200*.07 = 35.95 25.95 + 200*.05 = 35.95

OpenStudy (anonymous):

I should have written this first: You are solving for CellPlus = CellBest CellPlus = 21.95 + .07m CellBest = 25.95 + .05m Put the 2 equal to each other and solve

OpenStudy (anonymous):

21.95 + 0.07m = 25.95 + 0.05m